# A Decision-Language Model (DLM) for Dynamic Restless Multi-Armed Bandit Tasks in Public Health

### 22 Feb 2024 v1

@@ 3 Background
Figure 1: Overview of the proposed DLM loop. We begin by providing three key context descriptions to the LLM. First, we provide a language command; we include the specific commands used in Appendix Table 1. We also provide arm feature context, which contains a list of relevant per-arm features available to use in the proposed reward, as well as syntax cues and the relevant subset of the codebase so that the proposed LLM reward functions may be used directly in code. From this context, the LLM then proposes candidate reward functions which are used to train optimal policies under proposed reward. These trained policies are then used to generate policy outcome comparisons, which show the distribution of rewards across key feature groups in simulation. Finally, we query an LLM to perform self-reflection Shinn et al. (2024); Ma et al. (2024) by choosing the best candidate reward aligning with the original language command; the selected best rewards are used as context for future reward generation iterations.
@@ 4 Decision-Language Model for RMABs
Below, we provide an overview of the proposed DLM reward generation technique. An overview of the proposed technique is provided in Fig. 1. We first describe the problem setting of reward generation, and motivate the use of LLMs for reward generation from a theoretical perspective. We then discuss the context is provided to the LLM for initial reward candidate generation. Then, we discuss the independent simulation stage which identifies an optimal policy $\pi^{R*}$ under a proposed reward $R$, and the subsequent policy outcome comparison stage that simulates the respective policy reward outcomes. To be clear, $\pi^{R*}$ references the optimal policy induced by a given reward function $R$. This does not refer to an optimal reward $*R*$. Finally, we discuss the LLM reflection stage which considers the policy outcome comparison and identifies successful candidates to guide future in-context learning for LLM reward generation.
@@ 4.1 Problem Setting: Reward Generation for RMABs
$\pi^{R*}=$ $\mathop{\mathrm{argmax}}_{\pi\in\Pi}\left\{\sum_{n=1}^{N}R\left(\bm{s}_{n},\bm{z}\right)+\beta\,\mathbb{E}\left[J\left(\bm{s}^{\prime}\right)\mid\bm{s},\pi\right]\right\}$
@@
In summary, our key objective is, given a human-specified language prompt $\ell$, to generate the reward function $R$ that maximizes base function $F$ through optimal policy $\pi^{R*}$ for reward $R$. We propose using LLMs to automatically propose this reward function $R$, and automatically refine proposed rewards through an iterative self-reflection procedure.
@@ 4.2 Sample Inefficiency of Fitness Function
In order to evaluate each task description that we provide, we design a ‘ground truth fitness function’. This is the reward function (assigning a reward to each state and action) against which the learned policies are evaluated. However, this fitness function is unknown to the LLM. The ‘base’ policy in our comparison is obtained by training an RL policy to maximize the discounted reward obtained by this ground truth fitness function and serves as a baseline to our experiments. We want to note that, despite being the ‘true reward function’, this might not be the best reward function to learn the optimal policy in a learning theoretic sense. This is usually the case when rewards are sparse (i.e, hard to obtain) from simple policies and the sparsity of the reward signal makes learning hard. Thus, reward shaping/ reward engineering has been extensively studied in the literature (Dann et al. (2023); Sutton and Barto (2018); Ng et al. (1999); Singh et al. (2010)). The idea here is to use domain knowledge and design a different, dense reward whose optimal policy is the same as that of the original reward. However, the learning algorithm deployed finds it much easier to learn the optimal policy via the shaped reward. While the existing works (eg Ng et al. (1999)) show that reward shaping methods can provably help learning in the single agent setting, their results apply to the multi-agent problem of RMABs, where the Lagrangian relaxation decouples the value function across arms and allows us to focus on separate single-agent MDPs. In this work, we eschew the careful human design of the reward function, and allow an LLM to assume this role. Our thesis is that the world knowledge the LLM obtains during its training, along with reflection, equips it with the domain knowledge essential to design efficient reward functions. Automated design of such reward functions have been explored in the literature: Mataric (1994); Sorg et al. (2010); Zheng et al. (2020). Here, a meta-network is trained to learn the best rewards along with the policy network which is trained to learn the optimal policy for a given reward. In our work, reflection serves a similar purpose as training of the meta-network but with the number of steps being dramatically smaller and the output being in the form of code, making it interpretable.
@@
8:     for reward $i=1$ to $K$ do
9:         Init. policy and critic nets. $\theta_{n},\phi_{n}\quad\forall n\in[N]$
@@
15:                  $a_{n}\sim\theta_{n}(s_{n},\lambda)\quad\forall n\in[N]$
@@
19:              Update each $(\theta_{n},\phi_{n})$ with PPO using $\mathcal{D}$
@@ 4.4 DRL Simulation
We evaluate each LLM-proposed reward function $R_{1:K}$ by training policy networks $\theta^{i}_{n}$ under each proposed reward $R_{i}$ for each arm in simulation $\forall n\in[N]$ (shown Alg. 1, lines 8:20). We consider a simulation space which defines the number of arms $N$, each with fixed, but hidden, transition dynamics. Following the procedure defined in Alg. 1, for a given reward $R_{i}$, we first sample an action-charge $\lambda$, which is used to decouple the learning of arm policies (Eq. 3, Alg. 1 line 13). Then, we simulate $n_{s}$ timesteps, sampling trajectories for each arm under the action-charge $\lambda$ (Alg. 1 lines 14:16). These trajectories are used in a buffer $\mathcal{D}$ (Alg. 1 line 17) to eventually update the policy and critic networks for each arm $(\theta_{n},\phi_{n})$, as proposed in Killian et al. (2022) (Alg. 1 lines 19:20). We compute advantage estimates for the actor from estimated Q-values.
Figure 2: Main results. We report interquartile mean Agarwal et al. (2021) and standard error. Task colors denote difficulty. Green tasks consider only one feature group (i.e. age, income, etc). Yellow tasks tasks consider more than one feature group, or single features not explicitly defined (e.g. Middle Income). Red tasks are highly ambiguous and require completely inferring the correct features (e.g. Early and Late Workers has no direct feature, requiring inference from preferred call times). Our method is able to match base reward performance in a completely automated fashion. We evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $500$ samples of $10$ steps each per task. For each RMAB, we have a total of $48$ arms with a budget of $16$.
@@ 4.5 Reflection Stage
We propose the reflection comparison stage to enable LLM self-reflection and iteration on proposed rewards. At a high level, the outcome comparison produces a string showing, under each proposed reward $R_{i}$, the final simulated distribution of reward across key feature groups. Distributions, for instance, may be shown over age, gender, income, or other pre-defined feature characteristic groupings. This comparison is used in place of a numerical, scalar feedback function. Prior works assume access to some base fitness functions for self-reflection ($F$ as defined above) Singh et al. (2009). In the resource-constrained public health domain, we do not assume we have access to this ground truth function and extend the reward input from a scalar to a distribution.
Thus, rather than providing feedback through the base fitness function, we instead show the distribution of states achieved across various feature groups $g\in\mathbf{G}$; for example, over age or income brackets. Intuitively, because we focus on resource allocation in the public health domain, we propose that simulated state distributions provide sufficient signal for LLM to analyze the efficacy of proposed rewards, and select top candidate examples for future in-context learning.
We provide an overview of the policy outcome comparison procedure in Alg. 2. Given trained policy and critic networks $\theta_{n},\phi_{n}$, final lambda value $\lambda$, and feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$, we first simulate over $n_{\text{s}}$ evaluation timesteps (Alg. 2 lines 7:11). Note that we store only accumulated agent states $\mathbf{S}$; we use accumulated per-agent states to later reflect on achieved allocation outcomes. We then compute the percentage of reward attributed to arms within each sub-group $g\in\mathbf{G}$ (Alg. 2 lines 13:16). Finally, we report final reward distributions in an output string. We then query an LLM to select the best candidate example given the original prompt $\ell$ and the outcome distribution string $\mathcal{Y}$ (Alg. 1 lines 22:24). This selected example is then added as context in the next iteration of reward generation (Alg. 1 line 25).
This approach has two key strengths. First, we enable greater transparency of our approach, by using reward distributions to evaluate the effectiveness of a policy under a specified reward $\pi^{R}$. By providing only simulated outcome distributions from simulation to the reflection LLM, we ensure that self-reflection focuses only on alignment between the stated language goal and the resource allocation outcomes of our public health settings, without assuming any access to ground truth reward. Second, we allow for greater flexibility; by investigating only the outcome distributions of the proposed rewards, rather than original policies themselves, we enable a model-agnostic approach to self-iteration in reward design.
@@ 5 Experimental Evaluation
1:Input: Trained policy, critic nets. $\theta_{n},\phi_{n}\forall n\in[N]$, action-charge $\lambda$, feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$
@@
8:     $a_{n}\sim\theta_{n}(s_{n},\lambda)\quad\forall n\in[N]$
@@ 5.1 Tasks, Baselines, and Metrics
We provide experimental results for 11 distinct resource allocation prompts, shown in Appendix, Table 1. We separate tasks into three broad difficulty levels: simple single-feature tasks, middle-complexity multiple-feature tasks, and challenging, highly ambiguous tasks requiring complete inference of relevant features (see Figure 2). For each prompt, we compute normalized reward over 10 seeds; we describe the normalization equation below. To evaluate our proposed Decision-Language Model (DLM), we compare against several baseline reward functions:
Random: This policy samples arms uniformly at random. Through normalization, we set random performance score to be 0, as our floor. This would be equivalent to a maximum entropy reward, or simply a constant non-negative reward for any action. No Action: This policy samples no arms at each timestep. In most cases this will yield a negative normalized score, although in certain sparse reward tasks this may be close to the random policy. Default: This policy is trained with the same underlying algorithm as DLM, but the reward function is set to prioritize all beneficiaries equally; thus, for each arm, reward=1 for a positive state and reward=0 for a negative state. Most prior work in RMABs assume this reward. Base: This baseline sets the reward equal to the ground truth fitness function that all other methods are evaluated against. We use this reward as the consistent standard against which we measure the performance of our proposed method and other baselines. To be clear, each method uses a different reward at train time, but uses Base reward at evaluation time. We normalize with Base so that normalized reward is 1 for each task. DLM(No Reflection): This version of our proposed method ablates the reflection and iterative refinement process. It allows us to assess the importance of the reflection loop in improving the quality of the LLM-generated reward functions. DLM(Reflection): This is the full version of our proposed method.
@@ 5.2 Training Details
We use the Gemini Pro LLM model Team et al. (2023) from Google to generate reward functions, and train our downstream policy with RL using PPO Schulman et al. (2017) and the generated reward function. The downstream policy is trained for $2$ epochs total to prevent overfitting. We do a total of $2$ rounds of reflection and generate $2$ candidate reward functions per round. As detailed in Figure 2, we evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $5000$ total steps per task.
@@ 5.3 DLM Success and Failure Cases
We present the experimental evaluation of DLM in Figure 2. We find that the proposed DLM technique is a capable reward designer across a broad range of tasks in public health, demonstrating, to the best of our knowledge, first example of LLMs as capable policy designers for resource allocation tasks in public health. Note that we evaluate under a sample-limited regime; thus, results demonstrate the ability to converge under limited samples to an optimal policy for a given prompt. Given this setting, we make several observations on the results in Figure 2.
@@
Benefits of reflection: We observe that the proposed DLM reflection technique, which grounds LLM reward generation by incorporating feedback from RMAB simulations, is able to consistently align policies closer to desired human preferences. We note, however, that the degree of improvement from reflection may vary by task type. For instance, in cases with extremely simple prompts, for example single feature group preferences, reward reflection may not substantially enhance allocation outcomes over the zeroshot LLM-proposed rewards (i.e. Age Distribution Tail Emphasis above).
Additionally, in highly ambiguous cases, such as Technologically Challenged, we find that reflection may not have sufficient prompt guidance to meaningfully iterate on proposed rewards. Further, for cases with extreme ambiguity, requiring high degrees of reasoning, we found that the LLM response would frequently state: ”There was not enough information specified to solve the task”; in these cases, no zeroshot reward is proposed, obviating the benefits of reflection. One potential solution, we suggest, may be to add a human in the proposal loop, to further specify prompts and refine the reward function in highly ambiguous cases. We leave these potential strategies for addressing highly ambiguous tasks to future work.
@@
Above, the LLM incorrectly uses features in the reward function, as an arm cannot be in both the second-oldest and the oldest age group simultaneously. Thus, in cases where improperly designed rewards are proposed, reflection may be helpful (see Appendix E). For more examples of generated reward functions, please see Appendix F.
Middle-complexity tasks: We observe that the reflection procedure may be most consistently helpful for the “middle range” of tasks that involve a composition of feature groups within the reasoning capabilities of the LLM. This may be particularly helpful in cases with complex goals that cannot easily be translated into updated policies by public health practitioners. For example, in ARMMAN’s Kilkari program (a national level program in India, or operational across 19 states in India) Bashingwa et al. (2021), different regions may have slightly different policy requirements, which a state program manager can achieve directly through language interfacing with DLM.

### 23 Feb 2024 v2

@@ 3 Background
Figure 1: Overview of the proposed DLM loop. We begin by providing three key context descriptions to the LLM. First, we provide a language command; we include all specific commands in Appendix E. We also provide arm feature context, which contains a list of relevant per-arm features available to use in the proposed reward, as well as syntax cues and the relevant subset of the codebase so that the proposed LLM reward functions may be used directly in code. From this context, the LLM then proposes candidate reward functions which are used to train optimal policies under proposed reward. These trained policies are then used to generate policy outcome comparisons, which show the distribution of states across key feature groups in simulation. Finally, we query an LLM to perform self-reflection Shinn et al. (2024); Ma et al. (2024) by choosing the best candidate reward aligning with the original language command; the selected best rewards are used as context for future reward generation iterations.
@@ 4 Decision-Language Model for RMABs
Below, we provide an overview of the proposed DLM reward generation technique (shown Fig. 1). We first describe the problem setting of reward generation, and motivate the use of LLMs for reward generation from a theoretical perspective. We then discuss the prompt provided to the LLM for initial reward candidate generation. Then, we discuss the independent simulation stage which identifies an optimal policy $\pi^{R*}$ under a proposed reward $R$, and the subsequent policy outcome comparison stage that simulates the respective policy reward outcomes. To be clear, $\pi^{R*}$ references the optimal policy induced by a reward function $R$. This does not refer to policy learned under an optimal reward $R^{*}$. Finally, we discuss the LLM reflection stage which performs policy outcome comparison and identifies successful candidates to guide future in-context learning for LLM reward generation.
@@ 4.1 Problem Setting: Reward Generation for RMABs
$\pi^{R*}=$ $\mathop{\mathrm{argmax}}_{\pi\in\Pi}\left\{\sum_{n=1}^{N}R\left(\bm{s}_{n},\mathbf{z}\right)+\beta\,\mathbb{E}\left[J\left(\bm{s}^{\prime}\right)\mid\bm{s},\pi\right]\right\}$
@@
In summary, our key objective is, given a human-specified language prompt $\ell$, to generate the reward function $R$ that maximizes base function $F$ through optimal policy $\pi^{R*}$ for reward $R$. We propose using LLMs to both automatically propose this reward function $R$, and automatically refine proposed rewards through an iterative self-reflection procedure.
@@ 4.2 Sample Inefficiency of Fitness Function
In order to evaluate each task description that we provide, we manually define a ‘ground truth fitness function’. This is the reward function against which the learned policies are evaluated. However, this fitness function is unknown to the LLM. The ‘base’ policy in our comparison is obtained by training an RL policy to maximize the discounted cumulative reward obtained by this ground truth fitness function and serves as a baseline to our experiments. We want to note that, despite being the ‘true reward function’, this might not be the best reward function to learn the optimal policy in a learning theoretic sense. This is usually the case when rewards are sparse (i.e, hard to obtain) from simple policies and the sparsity of the reward signal makes learning hard. Thus, reward shaping/ reward engineering has been extensively studied in the literature (Dann et al. (2023); Sutton and Barto (2018); Ng et al. (1999); Singh et al. (2010)). The idea here is to use domain knowledge and design a different, dense reward whose optimal policy is the same as that of the original reward. However, the learning algorithm deployed finds it much easier to learn the optimal policy via the shaped reward. While the existing works (eg Ng et al. (1999)) show that reward shaping methods can provably help learning in the single agent setting, their results apply to the multi-agent problem of RMABs, where the Lagrangian relaxation decouples the value function across arms and allows us to focus on separate single-agent MDPs. In this work, we eschew the careful human design of the reward function, and allow an LLM to assume this role. Our thesis is that the world knowledge the LLM obtains during its training, along with reflection, equips it with the domain knowledge essential to design efficient reward functions. Automated design of such reward functions have been explored in the literature: Mataric (1994); Sorg et al. (2010); Zheng et al. (2020). Here, a meta-network is trained to learn the best rewards along with the policy network which is trained to learn the optimal policy for a given reward. In our work, reflection serves a similar purpose as training of the meta-network but with the number of steps being dramatically smaller and the output being in the form of code, making it interpretable.
@@
8:     Init. policy and critic net. $\theta,\phi$
9:     for reward $i=1$ to $K$ do
@@
15:                  $a_{n}\sim\theta(s_{n},\lambda)\quad\forall n\in[N]$
@@
19:              Update $(\theta,\phi)$ with PPO using $\mathcal{D}$
@@ 4.4 DRL Simulation
We evaluate each LLM-proposed reward function $R_{1:K}$ by training a policy network $\theta^{i}$ under each proposed reward $R_{i}$ (shown Alg. 1, lines 8:20). We consider a simulation space which defines the number of arms $N$, each with fixed, but hidden, transition dynamics. Following the procedure defined in Alg. 1, for a given reward $R_{i}$, we first sample an action-charge $\lambda$, which is used to decouple the learning of arm policies (Eq. 3, Alg. 1 line 13). Then, we simulate $n_{s}$ timesteps, sampling trajectories for each arm under the action-charge $\lambda$ (Alg. 1 lines 14:16). These trajectories are used in a buffer $\mathcal{D}$ (Alg. 1 line 17) to eventually update the policy and critic network$(\theta,\phi)$, similar to Killian et al. (2022) (Alg. 1 lines 19:20). We compute advantage estimates for the actor from estimated Q-values. Note we only learn a single policy/critic network for computational efficiency as opposed to Killian et al. (2022) which learns a network for each arm.
Figure 2: Main results. We report the interquartile mean of the normalized performance Agarwal et al. (2021) and standard error. Task colors denote difficulty. Green tasks consider only one feature group (i.e. age, income, etc). Yellow tasks consider more than one feature group, or single features not explicitly defined (e.g. Middle Income). Red tasks are highly ambiguous and require completely inferring the correct features (e.g. Early and Late Workers has no direct feature in the dataset, requiring inference from known features of preferred call times). Our method is able to match base reward performance for most tasks in a completely automated fashion. We evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $500$ samples of $10$ steps each per task. For each RMAB, we have a total of $48$ arms with a budget of $16$.
@@ 4.5 Reflection Stage
We propose the reflection comparison stage to enable LLM self-reflection and iteration on proposed rewards. The outcome comparison produces language output showing, under each proposed reward $R_{i}$, the final simulated distribution of states across key feature groups. Distributions, for instance, may be shown over age, gender, income, or other pre-defined feature groupings. This comparison is used in place of a numerical, scalar feedback function. Prior works Ma et al. (2024) assume access to some base fitness functions for self-reflection ($F$ as defined above) Singh et al. (2009) and also do not consider the multi-agent setting. In the resource-constrained public health domain, we do not assume we have access to this ground truth function. Thus, rather than providing feedback from the base fitness function, we instead show state-feature distributions, or the states achieved across feature groups $g\in\mathbf{G}$; for example, over age or income brackets. Intuitively, because we focus on resource allocation in the public health domain, we propose that simulated state-feature distributions provide sufficient signal for the LLM to analyze the efficacy of proposed rewards, and select top candidate examples for future in-context learning.
We provide an overview of the policy outcome comparison procedure in Alg. 2. Given trained policy and critic network $\theta,\phi$, final lambda value $\lambda$, and feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$, we first simulate over $n_{\text{s}}$ evaluation timesteps (Alg. 2 lines 7:11). Note that we store only accumulated agent states $\mathbf{S}$; we use accumulated per-agent states to later reflect on achieved allocation outcomes. We then compute the percentage of reward attributed to arms within each sub-group $g\in\mathbf{G}$ (Alg. 2 lines 13:16). Finally, we report final state-feature distributions in an output string. We then query an LLM to select the best candidate example given the original prompt $\ell$ and the outcome distribution string $\mathcal{Y}$ (Alg. 1 lines 22:24). This selected example is then added as context in the prompt of the next iteration of reward generation (Alg. 1 line 25).
This approach has two key strengths. First, we enable greater transparency of our approach, by using state-feature distributions to evaluate the effectiveness of a policy under a specified reward $\pi^{R*}$. By providing only simulated outcome distributions to the reflection LLM, we ensure that self-reflection focuses only on alignment between the stated language goal and the resource allocation outcomes of our public health settings, without assuming any access to ground truth reward. Second, we allow for greater flexibility; by investigating only the outcome distributions of the proposed rewards, rather than original policies themselves, we enable a model-agnostic approach to self-iteration in reward design.
@@ 5 Experimental Evaluation
1:Input: Trained policy, critic net. $\theta,\phi$, action-charge $\lambda$, feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$
@@
8:     $a_{n}\sim\theta(s_{n},\lambda)\quad\forall n\in[N]$
@@ 5.1 Tasks, Baselines, and Metrics
We provide experimental results for 11 distinct resource allocation prompts, given in Appendix E. We separate tasks into three broad difficulty levels: simple single-feature tasks, middle-complexity multiple-feature tasks, and challenging, highly ambiguous tasks requiring complete inference of relevant features (see Figure 2). For each prompt, we compute normalized performance over 10 seeds; we describe the normalization equation below. To evaluate our proposed Decision-Language Model (DLM), we compare against several baseline reward functions:
Random: This policy samples arms uniformly at random. Through normalization, we set the random reward function performance score to be 0, as our floor. This would be equivalent to solely maximizing for entropy, or simply a constant non-negative reward for any action. No Action: This policy samples no arms at each timestep. In most cases this will yield a negative normalized score, although in certain sparse reward tasks this may be close to the random policy. Default: This policy is trained with the same underlying algorithm as DLM, but the reward function is set to prioritize all beneficiaries equally; thus, for each arm, reward=1 for a positive state and reward=0 for a negative state. Most prior work in RMABs assume this reward. Base: This baseline sets the reward equal to the ground truth fitness function that all other methods are evaluated against. We use this reward as the consistent standard against which we measure the performance of our proposed method and other baselines. To be clear, each method uses a different reward at train time, but uses Base reward at evaluation time. We normalize with Base so that normalized performance using Base is 1 for each task. DLM (No Reflection): This version of our proposed method ablates the reflection and iterative refinement process. It allows us to assess the importance of the reflection loop in improving the quality of the LLM-generated reward functions. DLM (Reflection): This is the full version of our proposed method.
@@ 5.2 Training Details
We use the Gemini Pro LLM model Team et al. (2023) from Google to generate reward functions, and train our downstream policy with RL using PPO Schulman et al. (2017) and the generated reward function. The downstream policy is trained for $2$ epochs total to simulate a few-sample regime and prevent overfitting. We do a total of $2$ rounds of reflection and generate $2$ candidate reward functions per round. As detailed in Figure 2, we evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $5000$ total steps per task.
@@ 5.3 DLM Success and Failure Cases
We present the experimental evaluation of DLM in Figure 2. We find that the proposed DLM technique is a capable reward designer across a broad range of tasks in public health. This demonstrates, to the best of our knowledge, first example of LLMs as capable policy designers for resource allocation tasks in public health. Note that we evaluate under a sample-limited regime; thus, results demonstrate the ability to converge under limited samples to an optimal policy for a given prompt. Given this setting, we make several observations on the results in Figure 2.
@@
Please see Appendix G for all generated reward functions.
Benefits of reflection: We observe that the proposed DLM reflection technique, which grounds LLM reward generation by incorporating feedback from RMAB simulations, is able to consistently align policies closer to the desired task. We note, however, that the degree of improvement from reflection may vary by task type. For instance, in cases with extremely simple prompts, for example single feature group preferences, reward reflection tends not to substantially enhance allocation outcomes over the zeroshot LLM-proposed rewards (i.e. Age Distribution Tail Emphasis above).
Additionally, for ambiguous prompts, such as Technologically Challenged, we find that reflection may not have sufficient guidance to meaningfully iterate on proposed rewards. Further, for cases with extreme ambiguity, requiring high degrees of reasoning, the LLM frequently responded similar to ”There was not enough information specified to solve the task”. Thus, no zeroshot reward was proposed, obviating the benefits of reflection. One potential solution may be to add a human in the proposal loop, to further specify prompts in highly ambiguous cases. We leave these potential strategies for addressing highly ambiguous tasks to future work.
@@
Above, the LLM incorrectly uses features in the reward function, as an arm cannot be in both the second-oldest and the oldest age group simultaneously. Thus, in cases where improperly designed rewards are proposed, reflection is often helpful (see Appendix F). For more examples of generated reward functions, please see Appendix G.
Middle-complexity tasks: We observe that the reflection procedure is most consistently helpful for the “middle range” of tasks that involve a composition of feature groups within the reasoning capabilities of the LLM. Many real world applications fall into this category. For example, in ARMMAN’s Kilkari program (a national level program in India operational across 19 states) Bashingwa et al. (2021), different regions may have slightly different policy requirements, which a state program manager can achieve directly through language interfacing with DLM. More generally, complex goals that cannot easily be translated into updated policies by public health practitioners also fall into this category.